The value of Kpat 25°C for the reaction 2NO (g) + Cl2 (g) +2NOCI(g) is 1.9 ~ 103atm-1. Calculate the value of Kc at the same temperature.
1
Expert's answer
2020-06-08T18:31:45-0400
"2NO(g)+Cl_2(g)\leftrightarrow2NOCl(g)"

In the reaction, (2+1) mol of gas reacts to give 2 mol of gas.

Hence, "\Delta n=2-3=-1."

Using, "K_p=K_c(RT)^{\Delta n}:"


"K_c=K_p(RT)^{-\Delta n}"

"K_c=1.9\times10^{-3}(0.08206(25+273))^1=4.646\times10^{-2}"